Calculate the volume (in mL) of 0.100 M CaCl2 needed to produce 1.00g of CaCO (s). There is an excess of Na2CO3


Molar mass of calcium carbonate= 100.09 g/mol


volume of calcium chloride =


1
Expert's answer
2021-03-22T07:57:07-0400

Through mole-mass relationship between CaCl2 and CaCO3,

1mole of CaCl2 produces 100.09g of CaCO3

Amount of CaCl2 produced 1g of CaCO3= (1×1/100.09) =0.01mole.g

Getting volume; n=CV where; n=number of mole,C=concentration and V=volume in dm^3 or liter.

V=n/C

V=0.01/0.1=0.1L.

= 0.1×1000=100mL.